OpenStudy (anonymous):

An original rectangle has a length of 12 and a width of 12. What happens to the area if the new width is three times the original width? Verify your answer.

OpenStudy (anonymous):

it will increase a=3(12)x12 a=432

OpenStudy (lalaly):

12x12=area 144=area when width is 3 times 12*3(12)=432 432/144=3 Therefore, the new area is 3 times of the original

OpenStudy (anonymous):

A=a.b A2=3a.b so the area will be 3 thime of the first area too